# Boosting Transferability in Vision-Language Attacks via Diversification along the Intersection Region of Adversarial Trajectory

### v1

@@ Appendix 0.C Detailed Results about Cross Model Transferability
We conduct experiments with a focus on assessing the transferability of adversarial examples across two widely adopted VLP model architectures: fused and aligned. The selected VLP models for evaluation include ALBEF, TCL, CLIP, and CLIP [13]. The chosen downstream vision-and-language (V+L) task for assessment is image-text retrieval [18]. Our methodology involves utilizing one of the four models to generate multimodal adversarial examples within the specified parameters of our proposed method. Subsequently, we aim to validate the effectiveness of the generated adversarial examples through a comprehensive set of experiments. These experiments encompass both self-attacks, where the model attacks itself, and attacks on the other three models. The evaluation framework includes a white-box attack and three distinct black-box attacks.
@@
In the image-text retrieval task, we employ key metrics—R@1, R@5, and R@10 to gauge the efficacy of our attacks. These metrics assess whether the adversarial examples fail to appear in the top-ranked positions during the retrieval process, serving as crucial indicators of attack success. Due to space constraints in the main text, we focus on presenting the R@1 metric for cross-model transferability experiments. In Table 5, we furnish a comprehensive overview of the R@1, R@5, and R@10 metrics specifically for the Flickr30k dataset. Our analysis delves into the performance of the source model and the transferability across models separately.
@@
When validating the effectiveness of our method in improving the transferability of cross VLP models on Flickr30K [31], we utilize a test set consisting of 1,000 images, with approximately 5 captions for each image. To further validate the reliability of our method more extensively, we also employ a larger MSCOCO [26] test set for evaluation. This set comprises 5,000 test images, with each image having approximately 5 captions as well. The attack settings on MSCOCO are identical to those on Flickr30K, and detailed experimental results can be found in Table 6.
@@
In the experiments conducted on the Flickr30K and MSCOCO datasets, we set the perturbation constraint to 8/255 and the single step size to 2/255, with a total of 10 steps. To further explore the effectiveness of our method under different perturbation settings, we compare it with existing state-of-the-art methods using another perturbation configuration. We utilize a reduced perturbation configuration as outlined in the SGA paper, with a constraint of 2/255, employing a single step size of 0.5/255, and conducting a total of 10 steps. The results can be seen in Table 7 and 8. Similar to the limitation at 8/255, our method achieves a slightly lower success rate on the source model compared to SGA, but still at a significantly high level. Furthermore, the transferability of adversarial examples generated by our method is much better than our baseline method, SGA. Our experimental results under different perturbations and datasets thoroughly demonstrate the effectiveness of our method in generating multimodal adversarial examples with higher transferability.
@@
| ALBEF | TCL | CLIP | CLIP | ||||||||||
@@
| ALBEF | TCL | CLIP | CLIP | ||||||||||
@@
| ALBEF | TCL | CLIP | CLIP | ||||||||||
@@
| ALBEF | TCL | CLIP | CLIP | ||||||||||
@@
| ALBEF | TCL | CLIP | CLIP | ||||||||||
@@
| ALBEF | TCL | CLIP | CLIP | ||||||||||
@@
| ALBEF | TCL | CLIP | CLIP | ||||||||||
@@
| ALBEF | TCL | CLIP | CLIP | ||||||||||

### v3

@@ Appendix 0.G Detailed Results about Cross Model Transferability
We conduct experiments with a focus on assessing the transferability of adversarial examples across two widely adopted VLP model architectures: fused and aligned. The selected VLP models for evaluation include ALBEF, TCL, CLIPViT, and CLIPCNN [20]. The chosen downstream vision-and-language (V+L) task for assessment is image-text retrieval [28]. Our methodology involves utilizing one of the four models to generate multimodal adversarial examples within the specified parameters of our proposed method. Subsequently, we aim to validate the effectiveness of the generated adversarial examples through a comprehensive set of experiments. These experiments encompass both self-attacks, where the model attacks itself, and attacks on the other three models. The evaluation framework includes a white-box attack and three black-box attacks.
@@
In the image-text retrieval task, we employ key metrics—R@1, R@5, and R@10 to gauge the efficacy of our attacks. These metrics assess whether the adversarial examples fail to appear in the top-ranked positions during the retrieval process, serving as crucial indicators of attack success. Due to space constraints in the main text, we focus on presenting the R@1 metric for cross-model transferability experiments. In Table A3, we furnish a comprehensive overview of the R@1, R@5, and R@10 metrics specifically for the Flickr30k dataset. Our analysis delves into the performance of the source model and the transferability across models separately.
@@
When validating the effectiveness of our method in improving the transferability of cross VLP models on Flickr30K [42], we utilize a test set consisting of 1,000 images, with approximately 5 captions for each image. To further validate the reliability of our method more extensively, we also employ a larger MSCOCO [36] test set for evaluation. This set comprises 5,000 test images, with each image having approximately 5 captions as well. The attack settings on MSCOCO are identical to those on Flickr30K, and detailed experimental results can be found in Table A4.
@@
In the experiments conducted on the Flickr30K and MSCOCO datasets, we set the perturbation constraint to 8/255 and the single step size to 2/255, with a total of 10 steps. To further explore the effectiveness of our method under different perturbation settings, we compare it with existing state-of-the-art methods using another perturbation configuration. We utilize a reduced perturbation configuration as outlined in the SGA paper, with a constraint of 2/255, employing a single step size of 0.5/255, and conducting a total of 10 steps. The results can be seen in Table A5 and A6. Similar to the limitation at 8/255, our method achieves a slightly lower success rate on the source model compared to SGA, but still at a significantly high level. Furthermore, the transferability of adversarial examples generated by our method is much better than our baseline method, SGA. Our experimental results under different perturbations and datasets thoroughly demonstrate the effectiveness of our method in generating multimodal adversarial examples with higher transferability.
@@
| ALBEF | TCL | CLIPViT | CLIPCNN | ||||||||||
@@
| ALBEF | TCL | CLIPViT | CLIPCNN | ||||||||||
@@
| ALBEF | TCL | CLIPViT | CLIPCNN | ||||||||||
@@
| ALBEF | TCL | CLIPViT | CLIPCNN | ||||||||||
@@
| ALBEF | TCL | CLIPViT | CLIPCNN | ||||||||||
@@
| ALBEF | TCL | CLIPViT | CLIPCNN | ||||||||||
@@
| ALBEF | TCL | CLIPViT | CLIPCNN | ||||||||||
@@
| ALBEF | TCL | CLIPViT | CLIPCNN | ||||||||||